6. Find the minimum sample size required in to construct an 85% confidence interval about the population proportion p of collPlease show work. Thank you.
Answer #1

Solution :

Given that,

\hat p = 0.112

1 - \hat p = 1 - 0.112 = 0.888

margin of error = E = 0.03

Z\alpha/2 = 1.44

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (1.44 / 0.03)2 * 0.112 * 0.888

= 230

sample size = 230